# New Primitive OBZCPs of Lengths up to 49

### v1

@@ III-E Sketch of the Proposed Algorithm
2:$MAX\_ACC\geq 1$, maximum correlation
3:$mid_{a,b}=(0|1)$, mid bit for $\mathbf{a}$ and $\mathbf{b}$ correspondingly
4:$lsb=(0|1)$, least significant bit for $\mathbf{a}$ and $\mathbf{b}$
5:$CHUNK\_SIZE\Leftarrow(N-3)/2$
6:$NUMBER\_OF\_CHUNKS\Leftarrow 1<<CHUNK\_SIZE$
7:for $\mathbf{c}\leftarrow 1$ to $NUMBER\_OF\_CHUNKS$ do
8:     for  $\mathbf{g}\leftarrow 0$ to $12^{CHUNK\_SIZE}-$ 1 do
9:         $\mathbf{a}\leftarrow(lsb,\mathbf{g},mid_{a},rev(\mathbf{g},\mathbf{c}),1)$
10:         Add $\mathbf{a}$ to $map[\rho(\mathbf{a})]$
11:     end for
12:     $\mathbf{b\_lower}\leftarrow(1,0,0,\ldots,0)$
13:     Calculate $\mathbf{b\_upper}$ such that Eq.(4) holds for $\mathbf{b}=(lsb,\mathbf{b\_lower},mid_{b},\mathbf{b\_upper},1)$
14:     for $j\leftarrow 1$ to $2^{CHUNK\_SIZE}$ do
15:         $\rho_{\mathbf{a}}(\tau)\leftarrow-\rho_{\mathbf{b}}(\tau),\tau=1,\ldots,(N-3)/2$
16:         for $\mathbf{a}\in map[\rho(\mathbf{a})]$ do
17:              if  $\mathbf{a}\geq\mathbf{b}$ then
18:                  Print $(\mathbf{a},\mathbf{b})$ if out-of-zone correlation $\leq MAX\_ACC$
19:              end if
20:         end for
21:         Generate next Gray code for $\mathbf{b\_lower}$ part by changing single bit $x$
22:         Update $\mathbf{b\_upper}$ part by changing bit $N-1-x$
@@
25:Filter out equivalent pairs
The proposed algorithm is based on one outer loop to iterate over all the chunks and two consecutive inner loops for 1) loading data in memory map and 2) iterating over Gray code by changing a single bit on each step and retrieving matching values from the map such that equation (4) holds.
Considering that the corresponding search in table between $\mathbf{a}$ and $\mathbf{b}$ has lookup complexity of $O(1)$, iterating over $2^{(N-3)/2}$ values and $2^{(N-3)/2}$ chunks results in total complexity of $O(2^{N})$. We execute this algorithm $8$ times with all possible combinations of $a_{(N-1)/2},b_{(N-1)/2},a_{0}$.
In the algorithm, $rev(\mathbf{g},\mathbf{c})$ is defined such that $\mathbf{g}\oplus rev(\mathbf{g},\mathbf{C})\in\mathbf{C}$.
Note that the most inner loop is usually of size $0,1,2,$ rarely up to $4$ and hence does not impact the overall complexity.
@@ IV Computer Search Results
It is noted that all the OBZCPs are presented in hexadecimal notation. This allows us to map every hexadecimal digit to four binary digits, i.e. ($000\rightarrow 0$ $0$, $1\rightarrow 0001$, …, $F\rightarrow 1111$) and by stripping initial zeros if necessary. For instance $24159FE$ 24 represents $1010110011111111000100100$.
@@
Column 1 Column 2 Column 3 Column 4
(144E4E10, 114693FA) (1EE4E4BA, 1BEC3950) (17CA7B3A, 14640A7C) (1FCADB38, 1C64AA7E)
(1D60D190, 129F5190) (15B04C28, 153183F6) (1F1AE682, 18AD64C0) (1A9FF192, 16CE3F2A)
(1835D190, 17CA5190) (1A9FDB38, 15605B38) (1F6CC570, 11D1AD20) (15C66FDA, 15C3E424)
(18D7E164, 1560232C) (1FCAF192, 16CE2A7E) (1D60FB3A, 146420D6) (13CDD580, 1185A836)
(1AEF4E13, 16F06EEB) (1045E4B9, 10446B47)
@@
Column 1 Column 2 Column 3 Column 4
(4ED3AE80, 43856426) (692FCE8C, 6479042A) (6EF9C5A5, 5D3703BD) (6C7FC945, 5EB32F1D)
(64AFE6B9, 640ACBC7) (779DA917, 44536F0F) (741985B7, 46D563EF) (76D0CC17, 46D563EF)
(5B4302B9, 4E054C13)
@@
Column 1 Column 2 Column 3 Column 4
(1EA6D8C0A, 1EA6C73F4) (140C726A0, 140C6D95E) (1935FF538, 1935E0AC6) (139F55F92, 139F4A06C)
(180CBE6AC, 180CA1952) (12A614C06, 12A60B3F8) (180B39EAC, 180B26152) (12A193406, 12A18CBF8)
(18CBFEAD8, 18CBE1526) (126154072, 12614BF8C) (18CBFEAD8, 136AF059C) (1F8B39ED4, 1F8B2612A)
(15219347E, 15218CB80) (1F8CBE6D4, 1F8CA192A) (152614C7E, 15260B380) (1C377DD68, 1C3762296)
(169DD77C2, 169DC883C) (1828D8272, 17D728272) (1DC97C288, 1DC963D76) (1763D6822, 1763C97DC)
@@
Since $N=35$ is the first odd length that does not have optimal pairs, we present on the left-hand-side of Fig. 1 on the magnitude plot of the AACF sums of the pair $(7905A9444,710C1A3B2)$. On the right-hand-side of Fig. 1, we show the magnitude plot of the optimal pair $(15BCD1FAF3340,\\ 10599EA0E984A)$ with length 49.
@@
V OBZCPs with Smallest Demerit factors
Finally, we leverage the Pursley-Sarwate criterion and present OBZCPs with smallest demerit factors for sequence lengths up to 49 in Table XIII.
TABLE XIII: OBZCPs with smallest demerit factors
@@ VI Conclusions
This paper has introduced a number of new primitive optimal and $Z$-optimal OBZCPs for $27\leq N\leq 49$ which are obtained through a computer search. By using selected algebraic properties of OBZCPs and with certain programming techniques, we have presented an algorithm with complexity of $O(2^{N})$. Next to the known OBZCP constructions for $N=33$ and $N=41$, we have found optimal sequence pairs for $N=37$ and $N=49$. Additionally, a list of OBZCPs with smallest demerit factors have been reported.

### v3

@@ III-E Sketch of the Proposed Algorithm
Figure 1: Structure of sequence $\mathbf{a}$ of length 7.
In the proposed algorithm, $\operatorname{rvr}(\mathbf{g},\mathbf{c})$ is defined as $\operatorname{rvr}(\mathbf{g}\oplus\mathbf{c})$, which is obtained by binary xor operation of $\mathbf{g}$ and $\mathbf{c}$, followed by reversing the resulting binary vector. Sequence $\mathbf{a}$ is constructed by $(\text{lsb},\mathbf{g},\text{mid}_{\mathbf{a}},\operatorname{rvr}(\mathbf{g},\mathbf{c}),1)$, where lsb and $\text{mid}_{\mathbf{a}}$ are single binary elements and $\mathbf{g}$, $\operatorname{rvr}(\mathbf{g},\mathbf{c})$ are binary vectors of length $(N-3)/2$. For illustration purpose, we show in Figure 1 a length-7 $\mathbf{a}$ which belongs to $\mathbf{C}$ for $\mathbf{g}=(1,1)$ and $\mathbf{c}=(1,0)$. In the algorithm, $\mathbf{b\_upper}$ and $\mathbf{b\_lower}$, two binary vectors of length $(N-3)/2$, are used for construction of $\mathbf{b}$.
@@
[…]
3:$\text{mid}_{\mathbf{a}}$, predefined value for $a_{(N-1)/2}$
4:$\text{mid}_{\mathbf{b}}$, predefined value for $b_{(N-1)/2}$
5:lsb, least significant bit of $\mathbf{a}$ and $\mathbf{b}$
6:for ${N_{\mathbf{c}}}\leftarrow 0$ to $12^{(N-3)/2}-$ 1 do
7:     for  $N_{\mathbf{g}}\leftarrow 0$ to $12^{(N-3)/2}-$ 1 do $\triangleright$ Add $\mathbf{a}\in\mathbf{C}$ to the map
8:         $\mathbf{a}\leftarrow(\text{lsb},\mathbf{g},\text{mid}_{\mathbf{a}},{\operatorname{rvr}(\mathbf{g},\mathbf{c})},1)$
9:         Add $\mathbf{a}$ to $\text{map}[\rho_{\mathbf{a}}]$
10:     end for
11:     $\mathbf{b\_lower}\leftarrow(1,0,0,\ldots,0)$triangleright Construct initial $\mathbf{b}$
12:     Calculate $\mathbf{b\_upper}$ such that Eq.(4) holds for $\mathbf{b}=(\text{lsb},\mathbf{b\_lower},\text{mid}_{\mathbf{b}},\mathbf{b\_upper},1)$ and given $\mathbf{c}$. $\mathbf{a}$’s bits are already aggregated with an xor in ${\mathbf{C}}$.
13:     for $j\leftarrow 1$ to $2^{(N-3)/2}$ do $\triangleright$ Iterate over possible $\mathbf{b}$ candidates via Gray code
14:         $\rho_{\mathbf{a}}(\tau)\leftarrow-\rho_{\mathbf{b}}(\tau),\tau=1,\ldots,(N-3)/2$
15:         for $\mathbf{a}\in\text{map}[\rho_{\mathbf{a}}]$ do
16:              if  $N_{\mathbf{a}}>N_{\mathbf{b}}$ then $\triangleright$ Skip duplication
[…]
18:              end if
19:         end for
20:         Generate next Gray code for $\mathbf{b\_lower}$  part by changing single bit $\mathbf{b}_{x}$
21:         Update $\mathbf{b\_upper}$ part by reverting single bit
22:$\mathbf{b}_{N-1-x}$
@@
25:Filter out the equivalent pairs
The proposed algorithm is based on one outer loop to iterate over all the sets and two consecutive inner loops for 1) loading data in the memory map and 2) iterating over the Gray code by changing a single bit on each step and retrieving matching values from the map such that equation (4) holds.
It is noted that the time complexity of the outer loops equals to the number of sets that we iterate on, i.e., $O(2^{(N-3)/2})$. The time complexity of the two consecutive inner loops is determined by the maximum of:
  • The complexity of the first inner loop which equals to the number of all constructed $\mathbf{a}$’s in each set, i.e., $O(2^{(N-3)/2)}$;
  • The complexity of the second inner loop which equals to the number of all the $\mathbf{b\_upper}$ values that we iterate on via Gray code, i.e., $O(2^{(N-3)/2})$. During the second inner loop, we construct all the possible $\mathbf{b}$ matching to the set $\mathbf{C}(\mathbf{c})$. It is also noted that the corresponding search in the table between $\mathbf{a}$ and $\mathbf{b}$ has lookup complexity of $O(1)$. Additionally, the most inner loop on line 10 is usually of size $0,1,2,$ and rarely up to $4$ and hence does not impact the overall complexity.
Thus, the total time complexity of the inner loops is $O(2^{(N-3)/2})$.
Based on the above analysis, the overall algorithm has a complexity of $O(2^{N})$ since it is executed 8 times with all possible combinations of $a_{(N-1)/2},b_{(N-1)/2},a_{0}$.
@@ IV Computer Search Results
It is noted that all the OBZCPs are presented in hexadecimal notation. This allows us to map every hexadecimal digit to four binary digits, i.e. ($000\rightarrow 0$ $0$, $1\rightarrow 0001$, …, $F\rightarrow 1111$) and by stripping initial zeros if necessary. For instance 159FE24 represents $1010110011111111000100100$.
@@
Column 1 Column 2 Column 3
(144E4E10, 114693FA) (17CA7B3A, 14640A7C) (1FCADB38, 1C64AA7E)
(1835D190, 17CA5190) (1A9FDB38, 15605B38) (1F6CC570, 11D1AD20)
(1AEF4E13, 16F06EEB)
@@
Column 1 Column 2 Column 3
(4ED3AE80, 43856426) (6C7FC945, 5EB32F1D) (64AFE6B9, 640ACBC7)
@@
Column 1 Column 2 Column 3
(1EA6D8C0A, 1EA6C73F4) (180CBE6AC, 180CA1952) (1F8B39ED4, 1F8B2612A)
@@
Since $N=35$ is the first odd length that does not have optimal pairs, we present on the left-hand-side of Fig. 1 on the magnitude plot of the AACF sums of the pair (7905A9444, 710C1A3B2). On the right-hand-side of Fig. 1, we show the magnitude plot of the optimal pair (15BCD1FAF3340,10599EA0E984A) with length 49.
@@
Our algorithm is implemented using the software language “Rust” known as blazingly fast and memory-efficient [25] and includes optimizations such as Rayon parallel computing library, usage of native CPU instructions via compilation flags, such as SIMD, loops unwinding, etc. Our source code can be found at https://github.com/peterkazakov/obzcp. The OBZCP search for lengths up to 41 was carried out over a 2.9 GHz 6-Core Intel i9 computer. The execution time over for $N=31$ and $N=33$ are $3$ and $14$ minutes, respectively. Each subsequent length requires approximately $4$ times more minutes, for instance $N=39$ requires a day of computations.
The other results for $N=43,45,47,49$ are found through Essex HPC.
V OBZCPs with Smallest PSC Values
Finally, we leverage the Pursley–Sarwate criterion and present OBZCPs with smallest PSC values for sequence lengths up to 49 in Table XIII.
TABLE XIII: OBZCPs with smallest PSC values
@@ VI Conclusions
This paper has introduced a number of new primitive optimal and $Z$-optimal OBZCPs for $27\leq N\leq 49$ which are obtained through a computer search. By using selected algebraic properties of OBZCPs and with certain programming techniques, we have presented an algorithm with time complexity of $O(2^{N})$. Next to the known OBZCP constructions for $N=33$ and $N=41$, we have found optimal sequence pairs for $N=37$ and $N=49$. Additionally, a list of OBZCPs with smallest PSC values have been reported.